# Can It Edit? Evaluating the Ability of Large Language Models to Follow Code Editing Instructions

### 05 Mar 2024 v4

@@ Appendix B Example Benchmark Items
In this section, we showcase four examples from the CanItEdit benchmark, which are representative of the types of problems present.
B.1. oop_refactor
Figure 7 details a task where the model refactors code using object-oriented programming (OOP) principles. Initially, the code is a function for formatting messages based on type. The refactoring involves creating TextMessage and ImageMessage as subclasses of an abstract Message class and implementing a MessageFactory for message construction.
This task provides an example of a revise edit (Section 3.1), focusing on reorganizing the code into an OOP style without adding new features. The transformation is quite significant, and the largest relative transformation in our dataset: from a single function to a multi-class OOP program.
The goal is to assess the model’s proficiency in converting functional code into well-structured OOP designs based on comprehensive instructions and for the model to restructure small programs into much larger ones. Our test suites verify both functional correctness and the proper hierarchical class structure.
B.2. group_theory
Figure 8 features a task to modify a class from representing group $C4$ to group $C8$, including its operations like inverse and product. This task is an exemplary revise edit, focusing on significantly adapting an existing class rather than adding new features.
@@
The problem also highlights domain-specific problems in CanItEdit, this one being set in the context of cyclic groups. Testing domain-specific edits is crucial, especially when comparing the capabilities of large proprietary models like GPT-4 with smaller open models. It requires the model to transform the C4 class (representing a 4-element cyclic group) into the C8 class (for an 8-element group), requiring extensive edits across various code sections. This complexity presents a significant test for other code editing approaches, such as fill-in-the-middle (Bavarian et al., 2022; Fried et al., 2023), which may struggle with multiple edit locations (Yee and Guha, 2023).
Key edits involve altering the size and elements methods. The necessary understanding for these modifications stems from group theory, which is not explicitly explained in the problem. This setup tests the model’s capability to execute domain-specific edits where contextual knowledge is implied rather than provided.
B.3. strategy
Figure 9 presents an open-ended problem where the model devises a game strategy to defeat the already implemented CornerStrategy in Tic Tac Toe. This task represents an evolve edit, focused on developing a new feature without altering existing classes.
@@
The uniqueness in this problem lies in the lack of providing rules for the game, but rather requiring the model to infer them through understanding of the code. Additionally, it leaves the strategy design entirely to the model’s discretion. Our tests ensure that the Game class remain intact and that the model’s strategy consistently outperforms CornerStrategy in the game.
B.4. sudoku_solver
@@
Descriptive Instruction Lazy Instruction
Abstract the code into an object-oriented version of itself. To do that, create an abstract class ‘Message(ABC)‘, which can be initialized with a ‘content‘ string. The class should have an abstract method ‘process(self)‘, which should return a string. Create two children classes ‘TextMessage‘ and ‘ImageMessage‘, which implement the ‘process‘ method. Finally, create a ‘MessageFactory‘ that has a static method ‘get_message(message_type, content) -¿ Message‘; static methods can be defined with the ‘@staticmethod‘ decorator. The ‘get_message‘ method should return ‘Message‘ corresponding to the ‘message_type‘ (either ‘text‘ or ‘image‘), and it should throw a ValueError if the ‘message_type‘ is not valid.
Make the code object-oriented. Specifically, create an abstract class ‘Message‘, and children classes ‘TextMessage‘ and ‘ImageMessage‘. The ‘Message‘ class should have a method ‘process(self)‘ that prints the message which was given to the constructor. Also, create a ‘MessageFactory‘ that has a static method ‘get_message(message_type, content) -¿ Message‘; should raise an exception if the message type is not supported.
Before Code Segment Reference After Solution
def process_message(message, message_type):
    if message_type == "text":
        return f"Processed text message: {message}"
    elif message_type == "image":
        return f"Processed image message with description:
        {message}"
@@
        return "Unknown message type"
Ω
from abc import ABC, abstractmethod
class Message(ABC):
    Abstract class for messages
    def __init__(self, content):
        self.content = content
    @abstractmethod
    def process(self):
        pass
class TextMessage(Message):
    Concrete class for TextMessage
    def process(self):
        return f"Processed text message: {self.content}"
class ImageMessage(Message):
    Concrete class for ImageMessage
    def process(self):
        return f"Processed image message with description: {self.content}"
class MessageFactory:
    Factory class for creating message objects
    @staticmethod
    def get_message(message_type, content):
        if message_type == "text":
            return TextMessage(content)
        elif message_type == "image":
            return ImageMessage(content)
        else:
            raise ValueError("Unknown message type")
Figure 7. The oop_refactor problem from CanItEdit. This is a prime example of a revise type of edit, as asks the model to refactor code using OOP principles, without adding any additional features.
Descriptive Instruction Lazy Instruction
@@
Edit the C4 class and its methods to represent the C8 group instead.
Before Code Segment Reference After Solution
import torch
import numpy as np
import torch.nn as nn
class C4(nn.Module):
    """Represents the C4 class of group theory, where each
    element represents a discrete rotation."""
    def __init__(self):
        super().__init__()
        self.register_buffer(’identity’, torch.Tensor([0.]))
    def size(self):
        """Outputs the size of this group."""
        return 4
    def elements(self):
        """Returns all the elements of this group"""
        return torch.tensor([0., np.pi / 2, np.pi, 3 * np.pi / 2])
    def product(self, h, g):
        """Compute the product of two elements g and h in the group C4"""
        return torch.remainder(h + g, 2 * np.pi)
    def inverse(self, h):
        """Computes the inverse of the element h in the group C4"""
        return torch.remainder(-h, 2 * np.pi)
    def matrix_representation(self, h):
        """Returns the matrix representation of this element"""
        cos_t = torch.cos(h)
        sin_t = torch.sin(h)
        representation = torch.tensor([
            [cos_t, -sin_t],
            [sin_t, cos_t]
        ], device=self.identity.device)
        return representation
Ω
import torch
import numpy as np
import torch.nn as nn
class C8(nn.Module):
    """Represents the C8 class of group theory, where each
    element represents a discrete rotation."""
    def __init__(self):
        super().__init__()
        self.register_buffer(’identity’, torch.Tensor([0.]))
    def size(self):
        """Outputs the size of this group."""
        return 8
    def elements(self):
        """Returns all the elements of this group"""
        delta = np.pi / 4
        return torch.tensor([0., delta, delta * 2, delta * 3,
                             delta * 4, delta * 5, delta * 6, delta * 7])
    def product(self, h, g):
        """Compute the product of two elements g and h in the group C8"""
        return torch.remainder(h + g, 2 * np.pi)
    def inverse(self, h):
        """Computes the inverse of the element h in the group C8"""
        return torch.remainder(-h, 2 * np.pi)
    def matrix_representation(self, h):
        """Returns the matrix representation of this element"""
        cos_t = torch.cos(h)
        sin_t = torch.sin(h)
        representation = torch.tensor([
            [cos_t, -sin_t],
            [sin_t, cos_t]
        ], device=self.identity.device)
        return representation
Figure 8. The group_theory problem from CanItEdit. This exemplifies the subset of domain-specific problems in our benchmark.
Descriptive Instruction Lazy Instruction
The following code describes a tic-tac-toe game which takes in two strategies and determines who wins if they play each other. The ‘Strategy‘ class defines an abstract method, ‘returnMove(board)‘, which returns a tuple representing where this strategy will move, given a board state. The ‘CornerStrategy‘ class is a subclass of ‘Strategy‘ with a concrete implementation of ‘returnMove(board)‘. The ‘Game‘ class constructor takes in two strategies. It has a method ‘player1Won‘ which determines if the first strategy provided will beat the other if they both take turns alternating between moves. There are two methods, ‘playerXWon‘ and ‘gameOver‘ which determine how a game is won and when it is over. Create a class ‘GoodStrategy‘ which extends ‘Strategy‘ such that ‘Game(GoodStrategy(), CornerStrategy()).player1Won()‘ returns ‘True‘. This can not be solved by modifying the ‘Game‘, ‘Strategy‘, or ‘CornerStrategy‘ classes in any way.
Create a strategy ‘GoodStrategy‘, that beats ‘CornerStrategy‘. Do not modify the ‘Game‘ class.
Before Code Segment After Code Segment
from abc import ABC
from abc import abstractmethod
from typing import List, Tuple
class Strategy(ABC):
    @abstractmethod
    def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
        ’’’Returns a tuple(row, column) which indicates where to move
           in a 3x3 grid.’’’
        pass
class CornerStrategy(Strategy):
    def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
        if board[0][0] == None: return (0, 0)
        elif board[0][2] == None: return (0, 2)
        elif board[2][0] == None: return (2, 0)
        elif board[2][2] == None: return (2, 2)
        else: raise Exception
class Game:
    def __init__(self, player1: Strategy, player2: Strategy):
        self.playerOne = player1
        self.playerTwo = player2
        self.board = [[None for _ in range(3)] for _ in range(3)]
    def player1Won(self):
        playerTurn = True
        while not self.playerXWon(True) and not self.playerXWon(False) and not self.gameOver():
             strat = self.playerOne if playerTurn else self.playerTwo
             move = strat.returnMove(self.board)
             self.board[move[0]][move[1]] = playerTurn
             playerTurn = not playerTurn
        if self.gameOver(): return False
        else: return self.playerXWon(True)
    def gameOver(self):
        for row in self.board:
            for col in row:
                if col == None: return False
        return True
    def playerXWon(self, x: bool):
        for i in range(3):
            if self.rowNX(i, x): return True
        for i in range(3):
            if self.colNX(i, x): return True
        downDiag = self.board[0][0] == x and self.board[1][1] == x and self.board[2][2] == x
        upDiag = self.board[2][0] == x and self.board[1][1] == x and self.board[0][2] == x
        return downDiag or upDiag
    def rowNX(self, n: int, x: bool):
        for col in self.board[n]:
            if col != x: return False
        return True
    def colNX(self, n: int, x: bool):
        for row in self.board:
            if row[n] != x: return False
        return True
Ω
from abc import ABC
from abc import abstractmethod
from typing import List, Tuple
class Strategy(ABC):
    @abstractmethod
    def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
        ’’’Returns a tuple(row, column) which indicates where to move
           in a 3x3 grid.’’’
        pass
class CornerStrategy(Strategy):
    def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
        if board[0][0] == None: return (0, 0)
        elif board[0][2] == None: return (0, 2)
        elif board[2][0] == None: return (2, 0)
        elif board[2][2] == None: return (2, 2)
        else: raise Exception
class GoodStrategy(Strategy):
    def __init__(self) -> None:
        super().__init__()
        self.turn = 0
    def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
        self.turn += 1
        if self.turn == 1: return (0, 1)
        elif self.turn == 2: return (1, 1)
        elif self.turn == 3: return (2, 1)
        raise Exception
class Game:
    def __init__(self, player1: Strategy, player2: Strategy):
        self.playerOne = player1
        self.playerTwo = player2
        self.board = [[None for _ in range(3)] for _ in range(3)]
    def player1Won(self):
        playerTurn = True
        while not self.playerXWon(True) and not self.playerXWon(False) and not self.gameOver():
             strat = self.playerOne if playerTurn else self.playerTwo
             move = strat.returnMove(self.board)
             self.board[move[0]][move[1]] = playerTurn
             playerTurn = not playerTurn
        if self.gameOver(): return False
        else: return self.playerXWon(True)
    def gameOver(self):
        for row in self.board:
            for col in row:
                if col == None: return False
        return True
    def playerXWon(self, x: bool):
        for i in range(3):
            if self.rowNX(i, x): return True
        for i in range(3):
            if self.colNX(i, x): return True
        downDiag = self.board[0][0] == x and self.board[1][1] == x and self.board[2][2] == x
        upDiag = self.board[2][0] == x and self.board[1][1] == x and self.board[0][2] == x
        return downDiag or upDiag
    def rowNX(self, n: int, x: bool):
        for col in self.board[n]:
            if col != x: return False
        return True
    def colNX(self, n: int, x: bool):
        for row in self.board:
            if row[n] != x: return False
        return True
Figure 9. The strategy problem from CanItEdit. This problem is a prime example of a evolve type of edit, and is characteristic in the open-endedness of the instructions, both descriptive and lazy.
Descriptive Instruction Lazy Instruction
This version of the sudoku solver and checker does not reflect the original game of sudoku; the original game also checks for the uniqueness of 3x3 subgrids in addition to the rows and columns. Update the ‘assert_uniq‘ function to add new constraints for all nine 3x3 subgrids, and update the ‘check_valid‘ function to make sure that input grids have unique 3x3 subgrids.

### 20 Mar 2024 v5

@@ Appendix C Example CanItEdit Benchmark Problems
We showcase four examples from the CanItEdit benchmark, which we believe are representative of the types of problems present in the dataset.
External Libraries in CanItEdit
Our benchmark includes 21 problems that import external libraries, which are libraries outside of Python’s standard environment. We report the list of external libraries used and their number of appearances in the dataset: NumPy (13), Pandas (6), SciPy (3), scikit-learn (3), PyTorch (3), Z3 (2), autograd (2), Flask (1), vLLM (1)
oop_refactor
Figure 8 details a task where the model refactors code using object-oriented programming (OOP) principles. Initially, the code is a function for formatting messages based on type. The refactoring involves creating TextMessage and ImageMessage as subclasses of an abstract Message class and implementing a MessageFactory for message construction. This task provides an example of a perfective edit, focusing on reorganizing the code into an OOP style without adding new features. The transformation is quite significant, and the largest relative transformation in our dataset: from a single function to a multi-class OOP program. The goal is to assess the model’s proficiency in converting functional code into well-structured OOP designs based on comprehensive instructions and for the model to restructure small programs into much larger ones. Our test suites verify both functional correctness and the proper hierarchical class structure.
group_theory
@@
Figure 9 features a task to modify a class from representing group $C4$ to group $C8$, including its operations like inverse and product. The problem highlights domain-specific problems in CanItEdit, this one being set in the context of cyclic groups. Testing domain-specific edits is crucial, especially when comparing the capabilities of large proprietary models like GPT-4 with smaller open models. It requires the model to transform the C4 class (representing a 4-element cyclic group) into the C8 class (for an 8-element group), requiring extensive edits across various code sections. This complexity presents a significant test for other code editing approaches, such as fill-in-the-middle (Bavarian et al., 2022; Fried et al., 2023), which may struggle with multiple edit locations (Yee & Guha, 2023). Key edits involve altering the size and elements methods. The necessary understanding for these modifications stems from group theory, which is not explicitly explained in the problem. This setup tests the model’s capability to execute domain-specific edits where contextual knowledge is implied rather than provided.
strategy
@@
Figure 10 presents an open-ended problem where the model devises a game strategy to defeat the already implemented CornerStrategy in Tic Tac Toe. This task represents an adaptive edit, focused on developing a new feature without altering existing classes. The uniqueness in this problem lies in the lack of providing rules for the game, but rather requiring the model to infer them through understanding of the code. Additionally, it leaves the strategy design entirely to the model’s discretion. Our tests ensure that the Game class remain intact and that the model’s strategy consistently outperforms CornerStrategy in the game.
sudoku_solver
@@
1def process_message(message, message_type):
2    if message_type == "text":
3        return f"Processed text message: {message}"
4    elif message_type == "image":
5        return f"Processed image message with description: {message}"
6    else:
7        return "Unknown message type"
(a) ‘before‘ code segment of the oop_refactor problem (Figure 8).
(b) Instructions for the oop_refactor problem (Figure 8).
Instruction Type
Abstract the code into an object-oriented version of itself. To do that, create an abstract class Message(ABC)‘, which can be initialized with a content string. The class should have an abstract method process(self)‘, which should return a string. Create two children classes TextMessage and ImageMessage‘, which implement the process method. Finally, create a MessageFactory that has a static method get_message(message_type, content) -> Message‘; static methods can be defined with the @staticmethod decorator. The get_message method should return Message corresponding to the message_type (either text or image‘), and it should throw a ValueError if the message_type is not valid.
Descriptive
Make the code object-oriented. Specifically, create an abstract class Message‘, and children classes TextMessage and ImageMessage‘. The Message class should have a method process(self)‘ that returns the message which was given to the constructor. Also, create a MessageFactory that has a static method get_message(message_type, content) -> Message‘; should raise an exception if the message type is not supported.
Lazy
[…]
        if message_type == text”:
[…]
        elif message_type == image”:
[…]
@@
[…]
(c) Canonical solution for the oop_refactor problem (Figure 8).
Figure 8: The oop_refactor problem from CanItEdit. This is a prime example of a perfective type of edit, as asks the model to refactor code using OOP principles, without adding any additional features.
1import torch
2import numpy as np
3import torch.nn as nn
4
5class C4(nn.Module):
6    """Represents the C4 class of group theory, where each
7      element represents a discrete rotation."""
8
9    def __init__(self):
10        super().__init__()
11        self.register_buffer(’identity’, torch.Tensor([0.]))
12
13    def size(self):
14        """Outputs the size of this group."""
15        return 4
16
17    def elements(self):
18        """Returns all the elements of this group"""
19        return torch.tensor([0., np.pi / 2, np.pi, 3 * np.pi / 2])
20
21    def product(self, h, g):
22        """Compute the product of two elements g and h in the group C4"""
23        return torch.remainder(h + g, 2 * np.pi)
24
25    def inverse(self, h):
26        """Computes the inverse of the element h in the group C4"""
27        return torch.remainder(-h, 2 * np.pi)
28
29    def matrix_representation(self, h):
30        """Returns the matrix representation of this element"""
31        cos_t = torch.cos(h)
32        sin_t = torch.sin(h)
33        representation = torch.tensor([
34            [cos_t, -sin_t],
35            [sin_t, cos_t]
36        ], device=self.identity.device)
37        return representation
(a) ‘before‘ code segment of the group_theory problem (Figure 9)
(b) Instructions for the group_theory problem (Figure 9).
Instruction Type
@@
Descriptive
Edit the C4 class and its methods to represent the C8 group instead
Lazy
1import torch
2import numpy as np
3import torch.nn as nn
4
5 AA   class C8(nn.Module):
6 AA       """Represents the C8 class of group theory, where each
7    element represents a discrete rotation."""
8
9    def __init__(self):
10        super().__init__()
11        self.register_buffer(’identity’, torch.Tensor([0.]))
12
13    def size(self):
14        """Outputs the size of this group."""
15 AAAA           return 8
16
17    def elements(self):
18        """Returns all the elements of this group"""
19 AAAA           delta = np.pi / 4
20 AAAA           return torch.tensor([0., delta, delta * 2, delta * 3,
21 AAAA                               delta * 4, delta * 5, delta * 6, delta * 7])
22
23    def product(self, h, g):
24 AAAA           """Compute the product of two elements g and h in the group C8"""
25        return torch.remainder(h + g, 2 * np.pi)
26
27    def inverse(self, h):
28 AAAA           """Computes the inverse of the element h in the group C8"""
29        return torch.remainder(-h, 2 * np.pi)
30
31    def matrix_representation(self, h):
32        """Returns the matrix representation of this element"""
33        cos_t = torch.cos(h)
34        sin_t = torch.sin(h)
35        representation = torch.tensor([
36            [cos_t, -sin_t],
37            [sin_t, cos_t]
38        ], device=self.identity.device)
39        return representation
(c) Canonical solution for the group_theory problem (Figure 9).
Figure 9: The group_theory problem from CanItEdit. This exemplifies the subset of domain-specific problems in our benchmark.
1from abc import ABC
2from abc import abstractmethod
3from typing import List, Tuple
4
5class Strategy(ABC):
6    @abstractmethod
7    def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
8        ’’’Returns a tuple(row, column) which indicates where to move
9          in a 3x3 grid.’’’
10        pass
11
12class CornerStrategy(Strategy):
13    def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
14        if board[0][0] == None: return (0, 0)
15        elif board[0][2] == None: return (0, 2)
16        elif board[2][0] == None: return (2, 0)
17        elif board[2][2] == None: return (2, 2)
18        else: raise Exception
19
20class Game:
21    def __init__(self, player1: Strategy, player2: Strategy):
22        self.playerOne = player1
23        self.playerTwo = player2
24        self.board = [[None for _ in range(3)] for _ in range(3)]
25
26    def player1Won(self):
27        playerTurn = True
28        while (not self.playerXWon(True) and not self.playerXWon(False)
29        and not self.gameOver()):
30            strat = self.playerOne if playerTurn else self.playerTwo
31            move = strat.returnMove(self.board)
32            self.board[move[0]][move[1]] = playerTurn
33            playerTurn = not playerTurn
34        if self.gameOver(): return False
35        else: return self.playerXWon(True)
36
37    def gameOver(self):
38        for row in self.board:
39            for col in row:
40                if col == None: return False
41        return True
42
43    def playerXWon(self, x: bool):
44        for i in range(3):
45            if self.rowNX(i, x): return True
46        for i in range(3):
47            if self.colNX(i, x): return True
48        downDiag = self.board[0][0] == x and self.board[1][1] == x and self.board[2][2] == x
49        upDiag = self.board[2][0] == x and self.board[1][1] == x and self.board[0][2] == x
50        return downDiag or upDiag
51
52    def rowNX(self, n: int, x: bool):
53        for col in self.board[n]:
54            if col != x: return False
55        return True
56
57    def colNX(self, n: int, x: bool):
58        for row in self.board:
59            if row[n] != x: return False
60        return True
(a) ‘before‘ code segment of the strategy problem (Figure 10).
(b) Instructions for the strategy problem (Figure 10).
Instruction Type
The following code describes a tic-tac-toe game which takes in two strategies and determines who wins if they play each other. The Strategy class defines an abstract method, returnMove(board)‘, which returns a tuple representing where this strategy will move, given a board state. The CornerStrategy class is a subclass of Strategy with a concrete implementation of returnMove(board)‘. The Game class constructor takes in two strategies. It has a method player1Won which determines if the first strategy provided will beat the other if they both take turns alternating between moves. There are two methods, playerXWon and gameOver which determine how a game is won and when it is over. Create a class GoodStrategy which extends Strategy such that Game(GoodStrategy(), CornerStrategy()).player1Won()‘ returns True‘. This can not be solved by modifying the Game‘, Strategy‘, or CornerStrategy classes in any way.
Descriptive
Create a strategy GoodStrategy‘, that beats CornerStrategy‘. Do not modify the Game class.
Lazy
1from abc import ABC
2from abc import abstractmethod
3from typing import List, Tuple
4
5class Strategy(ABC):
6    @abstractmethod
7    def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
8        ’’’Returns a tuple(row, column) which indicates where to move
9          in a 3x3 grid.’’’
10        pass
11
12class CornerStrategy(Strategy):
13    def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
14        if board[0][0] == None: return (0, 0)
15        elif board[0][2] == None: return (0, 2)
16        elif board[2][0] == None: return (2, 0)
17        elif board[2][2] == None: return (2, 2)
18        else: raise Exception
19
20 AAAA   class GoodStrategy(Strategy):
21 AAAA       def __init__(self) -> None:
22 AAAA           super().__init__()
23 AAAA           self.turn = 0
24 AAAA       def returnMove(self, board: List[List[bool]]) -> Tuple[int, int]:
25 AAAA           self.turn += 1
26 AAAA           if self.turn == 1: return (0, 1)
27 AAAA           elif self.turn == 2: return (1, 1)
28 AAAA           elif self.turn == 3: return (2, 1)
29 AAAA           raise Exception
30
31class Game:
32    def __init__(self, player1: Strategy, player2: Strategy):
33        self.playerOne = player1
34        self.playerTwo = player2
35        self.board = [[None for _ in range(3)] for _ in range(3)]
36    def player1Won(self):
37        ...
38    def gameOver(self):
39        ...
40    def playerXWon(self, x: bool):
41        ...
42    def rowNX(self, n: int, x: bool):
43        ...
44    def colNX(self, n: int, x: bool):
45        ...
(c) Canonical solution for the strategy problem (Figure 10).
Figure 10: The strategy problem from CanItEdit. This problem is a prime example of a adaptive type of edit, and is characteristic in the open-endedness of the instructions, both descriptive and lazy.
1from typing import List, Optional
2from z3 import ArithRef, Int, Solver, Distinct, And, sat, IntVal
3
4def make_9x9_z3_board(board_text: str, solver: Solver) -> List[List[ArithRef]]:
5    """
6    Creates a board of z3 variables from a string representation of a board.
7    For unknown cells, make the value be 0, and for known cells, make the value
8    be a number from 1-9.
9    """
10    board = []
11    for line_counter, line in enumerate(board_text.splitlines()):
12        row = []
13        for char_counter, character in enumerate(line.strip()):
14            if character.isdigit():
15                num = int(character)
16                # 0 is unknown
[…]
18                if num == 0:
19                    solver.add(And(cell >= 1, cell <= 9))
20                    row.append(cell)
21                elif 0 < num < 10:
22                    solver.add(cell == IntVal(num))
23                    row.append(cell)
24        if len(row) != 9:
25            raise ValueError(
26                f"Invalid column count of board, must be 9, got {len(row)}")
27        board.append(row)
28
29    if len(board) != 9:
30        raise ValueError(
31            f"Invalid row count of board, must be 9, got {len(board)}")
32
33    return board
34
35def assert_uniq(solver: Solver, z3_board: List[List[ArithRef]]):
36    # Assert rows unique
37    for row in z3_board:
38        solver.add(Distinct(row))
39
40    # Assert columns unique
41    for col in zip(*z3_board):
42        solver.add(Distinct(col))
43
44def print_board(board: List[List[int]]):
45    for row in board:
46        print(row)
47
48def check_valid(board: List[List[int]]) -> bool:
49    for row in board:
50        if len(set(row)) != 9:
51            return False
52
53    for col in zip(*board):
54        if len(set(col)) != 9:
55            return False
56
57    return True
58
59def solve(board_text: str) -> Optional[List[List[int]]]:
60    solver = Solver()
61    z3_board = make_9x9_z3_board(board_text, solver)
62    board: List[List[int]] = [[] for _ in range(9)]
63    assert_uniq(solver, z3_board)
64    if solver.check() == sat:
65        model = solver.model()
66        for i, row in enumerate(z3_board):
67            row = [model.evaluate(cell).as_long()  # type: ignore
68                  for cell in row]
69            board[i] = row
70        return board
71    else: return None
(a) ‘before‘ code segment of the sudoku_solver problem (Figure 11).
(b) Instructions for the sudoku_solver problem (Figure 11).
Instruction Type
 This version of the sudoku solver and checker does not reflect the original game of sudoku; the original game also checks for the uniqueness of 3x3 subgrids in addition to the rows and columns. Update the assert_uniq function to add new constraints for all nine 3x3 subgrids, and update the check_valid function to make sure that input grids have unique 3x3 subgrids.
Descriptive